OpenStudy (kimo_indie):

Which line is parallel to the line that passes through the points (1, 7) and (-3, 4)

OpenStudy (alexandervonhumboldt2):

the slope of parallel line is same. use slope formula (y1-y2)/(x1-x2)=(7-4)/(1+3)=3/4. the slope of given line is 3/4. thus the slope of line parallel to it is also 3/4. y=3/4x+b there is no given point though which the parallel line passes. so there are millions of equations: y=3/4 y=3/4+1 y=3/4+0.9 and so on

OpenStudy (alexandervonhumboldt2):

woops mistake y=3/4x y=3/4x+1 y=3/4x+0.9 and so on

OpenStudy (kimo_indie):

thnx!